OpenStudy (anonymous):

Solve x^4 = 36x^2. Identify any double roots. A. {0 (double root), 6, –6} B. {–3, 3, 2 (double root)} C. {0, 3, 2} D. {6, –6}

OpenStudy (anonymous):

Can anybody please explain how to solve this?

OpenStudy (anonymous):

@whpalmer4 @whydoihavetosignupl @EffImHighOffSkittles @eaileen @R00nnYBraiNsbiG @terenzreignz @uri @IloveYou* @primeralph @prettyboy20202 @ajprincess @AnElephant @SnuggieLad @snorl4x @Directrix @dan815 @divagirl421 @DarkKat @Dmitt @faix95 @GOODMAN @hobbs978 @jhonyy9 @Jack17 @KawasumiKimiko @Luigi0210 @Loujoelou @love_jessika15 @lkkk @Callisto @countonme123 @ChrisJP @nincompoop @mukushla @marylou004 @mlb004 @MissHN182 @thefon19 @ryanjnfish @ladycomical @DDCamp @ryanjnfish @jd1028 @phantomofthemathwork @jtrich13 @ajprincess @PaulaCamacho @laxman03 @beccabee77 @babyslapmafro @whpalmer4 @jtrich13 @resident2013 @sumi29 @R00nnYBraiNsbiG @tfuller @Frostbite @SnuggieLad @raque766 @physicsboffin @Falco276 @e.cociuba @blurbendy @kausarsalley @sakigirl @RANE @Danny_Boy @SomeGirl1999 @Jamierox4ev3r @iforgot @lilkittykat36 @n14r96 @ChanceBBrick @knowel @meldo16 @Helpmeplease99 @Manda_luvsu @baseballer20142 @jhalt @RIUKUS @harpelletjain @Math81 @Frank @WildKat @maurimath @shamim @AliaKnight @MayMay_69 @amyamanda @victoria650 @Social @Whatever1 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @Zoe_Bailey @RIUKUS @henendez @shaneetta @niah2411 @harpelletjain @Math81 @WildKat @eidan @Hassan07 @Asuman @Frank @bani2100 @turtlebuddy @maurimath @lizzy89 @dborder @peush @EricaaaaB @yehonatan @tori45 @Ojasmin @Markbruce @april115 @RH @shamim @Ryaan @emilyhaddad @Jamierox4ev3r @RANE @NoelGreco @shkrina @lakenclark @Fazeelayaz @sarahusher @JazzyFrizo16 @music101 @HaperFink22 @whydoihavetosignup1 @domu @ashleyjade @Gikundab @Will.Neko @QuinterrianW_23 @tillman--- @baseballer20142 @Manda_luvsu @Charollete @amyamanda @harpreetsk @Whatever1 @HaydenBurden1996 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @shaneetta

OpenStudy (anonymous):

A is the correct answer.

OpenStudy (anonymous):

@Charollete can you please explain? if you don't mind

OpenStudy (anonymous):

Sure: bring over x^4 so you'd have \[36x ^{2}-x ^{4}\] Factor an \[x ^{2} \] out so you'd have:\[x ^{2}(36-x ^{2})\] Set \[x ^{2} = 0\] and \[(36-x ^{2}) = 0\] You find that x = 0 in the \[x ^{2}\] portion. Factor the second portion: \[(36-x ^{2})\] You find that it's a difference of squares: (x-6) and (x+6) set both of those equal to 0: (x-6) = 0 and (x+6) = 0 x = 6 and x = -6 and x = 0

OpenStudy (anonymous):

Thank YOU SO MUCH!!!

OpenStudy (anonymous):

No problem =)